# Extended mean field control problems with constraints: The generalized Fritz-John conditions and Lagrangian method

### v1

@@ 4.2 Constrained (extended) mean field game
In this section, we discuss the case of C-(E)MFG problem. We interpret the constrained MFG problem as a constrained stochastic control problem with the fixed point consistency condition.
We formulate the C-(E)MFG as follows:
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\rho_{s},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\rho_{s},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\rho_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$,}\end{cases}$ (61)
where the measure flow $\rho=(\rho_{t})_{t\in[0,T]}$ takes values in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$. Here, $\widehat{\alpha}=(\widehat{\alpha}_{t})_{t\in[0,T]}$ is an (admissible) optimal control and $\widehat{X}=(\widehat{X}_{t})_{t\in[0,T]}$ is the resulting state process. The measure flow $\rho=(\rho_{t})_{t\in[0,T]}$ satisfying $\rho_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$ is called a mean field equilibrium (MFE) of the C-(E)MFG problem (61).
Corollary 4.2 (SMP for C-(E)MFG problem).
Let $\rho_{t}$ be an MFE to the C-(E)MFG problem (61). Then, there exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that
@@
    the positive radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for all $i\in I_{1}$,
@@
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},\rho_{s},Y_{s},Z_{s})\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ (63)
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$
@@
4.3 An example: dynamic constrained LQ-MFC problem
@@
under dynamic state-control constraint described by, $\mathbb{P}$-a.s.
@@
$\begin{cases}\displaystyle\mathrm{d}Y_{t}=-\left\{b_{1}Y_{t}+b_{2}\mathbb{E}[Y_{t}]+\sigma_{1}Z_{t}+\sigma_{2}\mathbb{E}[Z_{t}]+q\widehat{X}_{t}\right\}\mathrm{d}t+h_{t}\eta_{t}\mathrm{d}t+Z_{t}\mathrm{d}W_{t},\\[8.00003pt] \displaystyle Y_{T}=\ell\{\widehat{X}_{T}-\mathbb{E}[\widehat{X}_{T}]\}.\end{cases}$ (68)
@@
$\begin{cases}\displaystyle\mathrm{d}Y_{t}=-\left\{b_{1}Y_{t}+b_{2}\mathbb{E}[Y_{t}]+\sigma_{1}Z_{t}+\sigma_{2}\mathbb{E}[Z_{t}]+q\widehat{X}_{t}\right\}\mathrm{d}t+Z_{t}\mathrm{d}W_{t},\\[8.00003pt] \displaystyle Y_{T}=\ell\{\widehat{X}_{T}-\mathbb{E}[\widehat{X}_{T}]\}.\end{cases}$ (70)
@@
which coincides with the result for the extended MFC problem in the LQ framework as studied in [9] and an analytic solution can be derived by assuming that $Y_{t}$ takes the form $\beta_{t}X_{t}+\zeta_{t}\mathbb{E}[X_{t}]$ and solving the corresponding Riccati equation similar to [9].
Acknowledgements L. Bo and J. Wang are supported by Natural Science Basic Research Program of Shaanxi (No. 2023-JC-JQ-05), Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSZ010) and Fundamental Research Funds for the Central Universities (No. 20199235177). X. Yu is supported by the Hong Kong RGC General Research Fund (GRF) under grant no. 15306523 and grant no. 15211524.
@@ Appendix A Proofs of Auxiliary Results
$|g_{i}(t_{2})-g_{i}(t_{1})|$ $\leq\mathbb{E}[|\phi^{i}(t_{1},X_{t_{1}},\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{2}})|]$
$\leq\mathbb{E}[|\phi^{i}(t_{1},X_{t_{1}},\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})|]+\mathbb{E}[|\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{2}})|]$
$:=v_{1}+v_{2}.$
Note that $v_{1}$ converges to 00 as $|t_{1}-t_{2}|\to 0$ by DCT and the continuity of $\phi^{i}$ in $(t,x)$. Furthermore, by DCT, $\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]=\mathbb{E}[\int_{t_{1}}^{t_{2}}|\alpha_{s}|^{2}\mathrm{d}s]$ converges to 00 as $|t_{1}-t_{2}|\to 0$ since $\alpha\in\mathbb{L}_{\mathbb{F}}^{2}$. For the term $v_{2}$, we have that, as $|t_{1}-t_{2}|\to 0$,
$v_{2}$ $=\mathbb{E}[|\langle D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}}),\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\rangle_{\mathcal{X}}|]+o(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}]\|_{\mathcal{X}})$
$\leq\left(\mathbb{E}[\|D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}\left(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}+o(\left(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}).$
By using Assumption 1-(A2), $\mathbb{E}[\|D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})\|_{\mathcal{X}}^{2}]<\infty$, and hence $v_{2}\to 0$ as $|t_{1}-t_{2}|\to 0$ due to the fact $\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]=\mathbb{E}[\int_{t_{1}}^{t_{2}}|\alpha_{s}|^{2}\mathrm{d}s]$. ∎
@@
For any $Z\in\mathbb{L}_{\mathbb{F}}^{2}$, the classical SDE theory (c.f. Karatzas and Shreve [28]) gives that the following SDE has a unique strong solution $Y\in\mathbb{L}_{\mathbb{F}}^{2}$:
@@
where, for $t\in[0,T]$, $\overline{Z}_{t}$ stands for the r.v. $Z_{t}$ instead of the evaluation $Z_{t}(\omega)$ at $\omega\in\Omega$. Then, we have constructed a self-mapping $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$ by $G(Z)=Y$. Thus, the existence of solution reduces to showing that the mapping $G$ admits a fixed point. For $i=1,2$, let $Y^{i}=(Y_{t}^{i})_{t\in[0,T]}$ be the corresponding solution associated to $Z^{i}\in\mathbb{L}_{\mathbb{F}}^{2}$. Let $M=M(T)>0$ be a generic constant that may differ from line to line. The standard moment estimation under the assumption (28) on coefficients $(B,\Sigma)$ results in
$\mathbb{E}\left[|Y_{t}^{1}-Y_{t}^{2}|^{2}\right]\leq M\int_{0}^{t}\left\{\mathbb{E}\left[|Y_{s}^{1}-Y_{s}^{2}|^{2}\right]+\mathbb{E}\left[|Z_{s}^{1}-Z_{s}^{2}|^{2}\right]\right\}\mathrm{d}s,~{}~{}\forall t\in[0,T].$
Applying the Gronwall’s inequality, we obtain that
$\mathbb{E}\left[|Y_{t}^{1}-Y_{t}^{2}|^{2}\right]\leq M\int_{0}^{t}\mathbb{E}\left[|Z_{s}^{1}-Z_{s}^{2}|^{2}\right]\mathrm{d}s,~{}~{}\forall t\in[0,T],$ (72)
which yields the continuity of $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$.
@@
Under the assumption (28), it holds from (72) that, for all $n\in\mathbb{N}$,
@@
This implies that $|Y^{n+1}-Y^{n}|_{\mathbb{L}_{\mathbb{F}}^{2}}^{2}=O(\frac{1}{n!})$ as $n\to\infty$. Thus, $(Y_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $\mathbb{L}_{\mathbb{F}}^{2}$, and hence there exists a $Y\in\mathbb{L}_{\mathbb{F}}^{2}$ such that $Y^{n}\to Y$, as $n\to\infty$, in $\mathbb{L}_{\mathbb{F}}^{2}$. Letting $n\to\infty$, we can see that $Y\in\mathbb{L}_{\mathbb{F}}^{2}$ is a fixed point of the mapping $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$. By using (72) and the Gronwall’s inequality again, we conclude the uniqueness of the solution to the SDE (30). ∎

### v2

@@ 4.2 Constrained (extended) mean field game
In this section, we discuss the constrained MFG problems (C-MFG) with the dynamic expectation constraints and/or the dynamic state-control constraints. We regard the constrained MFG problem as a constrained stochastic control problem with the fixed point consistency condition.
We formulate the C-MFG as follows:
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[3.99994pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\rho_{s},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\rho_{s},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[6.00006pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\\[6.00006pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\rho_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$,}\end{cases}$ (45)
where the measure flow $\rho=(\rho_{t})_{t\in[0,T]}$ takes values in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$. Here, $\widehat{\alpha}=(\widehat{\alpha}_{t})_{t\in[0,T]}$ is an (admissible) optimal control and $\widehat{X}=(\widehat{X}_{t})_{t\in[0,T]}$ is the resulting state process. The measure flow $\rho=(\rho_{t})_{t\in[0,T]}$ satisfying $\rho_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$ is called a mean field equilibrium (MFE) of the C-MFG problem (45).
Corollary 4.2 (SMP for C-MFG problem).
Let $\rho_{t}$ be an MFE to the C-MFG problem (45). Then, there exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that
@@
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},\rho_{s},Y_{s},Z_{s})\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ (47)
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$
@@
4.3 An illustrative example: constrained LQ-MFC problem
@@
under dynamic state-law-control constraint described by, $\mathbb{P}$-a.s.
@@
$\mathrm{d}Y_{t}=-\{b_{1}Y_{t}+b_{2}\mathbb{E}[Y_{t}]+\sigma_{1}Z_{t}+\sigma_{2}\mathbb{E}[Z_{t}]+q\widehat{X}_{t}\}\mathrm{d}t+h_{t}\eta_{t}\mathrm{d}t+Z_{t}\mathrm{d}W_{t},~{}Y_{T}=\ell\{\widehat{X}_{T}-\mathbb{E}[\widehat{X}_{T}]\}.$ (52)
@@
$\mathrm{d}Y_{t}=-\{b_{1}Y_{t}+b_{2}\mathbb{E}[Y_{t}]+\sigma_{1}Z_{t}+\sigma_{2}\mathbb{E}[Z_{t}]+q\widehat{X}_{t}\}\mathrm{d}t+Z_{t}\mathrm{d}W_{t},~{}Y_{T}=\ell\{\widehat{X}_{T}-\mathbb{E}[\widehat{X}_{T}]\}.$ (54)
@@
which coincides with the result for the extended MFC problem in the LQ framework as studied in [9] and an analytic solution can be derived by assuming that $Y_{t}$ takes the form $\beta_{t}X_{t}+\zeta_{t}\mathbb{E}[X_{t}]$ and solving the associated Riccati equations as in [9].
Acknowledgements L. Bo and J. Wang are supported by National Natural Science of Foundation of China (No. 12471451), Natural Science Basic Research Program of Shaanxi (No. 2023-JC-JQ-05), Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSZ010). X. Yu is supported by the Hong Kong RGC General Research Fund (GRF) under grant no. 15306523 and grant no. 15211524.
@@ Appendix A Proofs of Auxiliary Results
Proof of Corollary 2.11.
In above, the equality can only be attained when $\lambda_{i}=0$ for $i\in I$. Thus, the equality (56) can be translated into $\sum_{j\in J}\mu_{j}\circ Dh_{j}(\widehat{x})=\xi\in N_{C}(\widehat{x})$. Here, $\xi=(\xi_{1},\xi_{2})\in\mathtt{X}_{1}^{*}\times\mathtt{X}_{2}^{*}$ as the dual space of a product space is the product of their dual spaces. Hence, we have $\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x})=\xi_{1}$. Moreover, by the definition of normal cone, we can conclude that by choosing $c=(c_{1},\widehat{x}_{2})\in C\subset\mathtt{X}_{1}\times\mathtt{X}_{2}$, where we recall $\widehat{x}=(\widehat{x}_{1},\widehat{x}_{2})\in C\subset\mathtt{X}$, we have
$\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x}_{1})=\xi_{1}=0.$ (58)
We argue by contradiction to show that $\lambda_{i}=0$ for $i\in I$. To this end, we first suppose that $\ker(\lambda_{1})^{c}\neq\varnothing$. Then, by the surjection in LICQ($\widehat{x}$), there is an $x_{0}\in\mathtt{X}_{1}$ s.t. $(D_{x_{1}}g_{i}(\widehat{x}_{1})(x_{0}),D_{x_{1}}h_{j}(\widehat{x}_{1})(x_{1}))_{i\in I,j\in J}$ belongs to the nonempty set $\ker(\lambda_{1})^{c}\times\prod_{i\in I\backslash\{1\}}\ker(\lambda_{i})\times\prod_{j\in J}\ker(\mu_{j})$, which yields that
@@
$|g_{i}(t_{2})-g_{i}(t_{1})|$ $\leq\mathbb{E}[|\phi^{i}(t_{1},X_{t_{1}},\mathcal{L}(X_{t_{1}}),\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\mathcal{L}(X_{t_{2}}),\widetilde{\alpha}^{t_{2}})|]$
$\leq\mathbb{E}[|\phi^{i}(t_{1},X_{t_{1}},\mathcal{L}(X_{t_{1}}),\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\mathcal{L}(X_{t_{2}}),\widetilde{\alpha}^{t_{1}})|]$
$\quad+\mathbb{E}[|\phi^{i}(t_{2},X_{t_{2}},\mathcal{L}(X_{t_{2}}),\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\mathcal{L}(X_{t_{2}}),\widetilde{\alpha}^{t_{2}})|]:=v_{1}+v_{2}.$
Note that $v_{1}$ converges to 00 as $|t_{1}-t_{2}|\to 0$ by DCT and the continuity of $\phi^{i}$ in $(t,x,\mu)$. Furthermore, by DCT, $\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]=\mathbb{E}[\int_{t_{1}}^{t_{2}}|\alpha_{s}|^{2}\mathrm{d}s]$ converges to 00 as $|t_{1}-t_{2}|\to 0$ since $\alpha\in\mathbb{L}_{\mathbb{F}}^{2}$. For the term $v_{2}$, we have that, as $|t_{1}-t_{2}|\to 0$,
$v_{2}$ $=\mathbb{E}[|\langle D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\mathcal{L}(X_{t_{2}}),\widetilde{\alpha}^{t_{1}}),\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\rangle_{\mathcal{X}}|]+o(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}]\|_{\mathcal{X}})$
$\leq\left(\mathbb{E}[\|D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\mathcal{L}(X_{t_{2}}),\widetilde{\alpha}^{t_{1}})\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}\left(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}+o(\left(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}).$
By using Assumption 1-(A2), $\mathbb{E}[\|D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\mathcal{L}(X_{t_{2}}),\widetilde{\alpha}^{t_{1}})\|_{\mathcal{X}}^{2}]<\infty$, and hence $v_{2}\to 0$ as $|t_{1}-t_{2}|\to 0$ due to the fact $\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]=\mathbb{E}[\int_{t_{1}}^{t_{2}}|\alpha_{s}|^{2}\mathrm{d}s]$. ∎
@@
For any $Z\in\mathbb{L}_{\mathbb{F}}^{2}$, the classical SDE theory (c.f. Karatzas and Shreve [29]) gives that the following SDE has a unique strong solution $Y\in\mathbb{L}_{\mathbb{F}}^{2}$ to
@@
where, for $t\in[0,T]$, $\overline{Z}_{t}$ stands for the r.v. $Z_{t}$ instead of the evaluation $Z_{t}(\omega)$ at $\omega\in\Omega$. Then, we have constructed a self-mapping $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$ by $G(Z)=Y$. Thus, the existence of solution reduces to showing that the mapping $G$ admits a fixed point. For $i=1,2$, let $Y^{i}=(Y_{t}^{i})_{t\in[0,T]}$ be the corresponding solution associated to $Z^{i}\in\mathbb{L}_{\mathbb{F}}^{2}$. Let $M=M(T)>0$ be a generic constant that may differ from line to line. The standard moment estimation under the assumption (21) on coefficients $(B,\Sigma)$ results in $\mathbb{E}\left[|Y_{t}^{1}-Y_{t}^{2}|^{2}\right]\leq M\int_{0}^{t}\left\{\mathbb{E}\left[|Y_{s}^{1}-Y_{s}^{2}|^{2}\right]+\mathbb{E}\left[|Z_{s}^{1}-Z_{s}^{2}|^{2}\right]\right\}\mathrm{d}s$ for $t\in[0,T]$. Applying the Gronwall’s lemma, we obtain $\mathbb{E}\left[|Y_{t}^{1}-Y_{t}^{2}|^{2}\right]\leq M\int_{0}^{t}\mathbb{E}\left[|Z_{s}^{1}-Z_{s}^{2}|^{2}\right]\mathrm{d}s$ for $t\in[0,T]$, which gives the continuity of $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$.
@@
Under Assumption (21), it holds that, for all $n\in\mathbb{N}$,
@@
This implies that $|Y^{n+1}-Y^{n}|_{\mathbb{L}_{\mathbb{F}}^{2}}^{2}=O(\frac{1}{n!})$ as $n\to\infty$. Thus, $(Y_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $\mathbb{L}_{\mathbb{F}}^{2}$, and hence there exists a $Y\in\mathbb{L}_{\mathbb{F}}^{2}$ such that $Y^{n}\to Y$, as $n\to\infty$, in $\mathbb{L}_{\mathbb{F}}^{2}$. Letting $n\to\infty$, we can see that $Y\in\mathbb{L}_{\mathbb{F}}^{2}$ is a fixed point of the mapping $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$. By Gronwall’s lemma again, we conclude the uniqueness of the solution to SDE (23). ∎